# Critic-Guided Reinforcement Unlearning in Text-to-Image Diffusion

### v2

@@ A.5 Advantage-weighted gradient form and Monte Carlo estimator
Since $B=0$, equation equation 19 reduces to the unbiased advantage-weighted gradient identity.
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{\tau\sim p_{\theta}(\cdot|c)}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)(r(x_{0},c)-b(x_{t}))\right]$
@@
$\widehat{\nabla_{\theta}J}=\frac{1}{N}\sum_{n=1}^{N}\sum_{t=1}^{T}\left[\nabla_{\theta}\log p_{\theta}(x_{t-1}^{(n)}|x_{t}^{(n)},c^{(n)})(r(x_{0}^{(n)},c^{(n)})-b(x_{t}^{(n)}))\right]$
@@
Thus choosing $b(x_{t})=V_{\phi}(x_{t})$, precisely the value function in equation 17, and training $V_{\phi}$ to regress the observed terminal reward $r(x_{0},c)$ (Monte Carlo regression) approximates the optimal least-squares baseline $b^{\star}(x_{t})=\mathbb{E}[r(x_{0},c)|x_{t}]$ and thereby reduces variance of the estimator without introducing bias.
Appendix B Training Algorithms
B.1 Critic Training Algorithm
Algorithm 1 Critic Training for CGRU
1:Dataset of prompts $\mathcal{D}_{c}=\{c_{i}\}_{i=1}^{N}$, diffusion model $p_{\theta}$, reward function $r(x_{0},c)$
2:Trained critic $V_{\phi}(x_{t},c,t)$
3:Initialize critic network $V_{\phi}$ with random parameters $\phi$
4:Initialize empty buffer $\mathcal{B}=\emptyset$
5:for epoch =1=1 to $E_{\text{critic}}$ do
6:  for each prompt $c\sim\mathcal{D}_{c}$ do
7:   Generate trajectory: $x_{T}\sim\mathcal{N}(0,I)$, $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t},c)$ for $t=T,\ldots,1$
8:   Compute terminal reward: $r_{\text{final}}=r(x_{0},c)$
9:   Store $(x_{t},c,t,r_{\text{final}})$ in $\mathcal{B}$ for all $t\in\{1,\ldots,T\}$
10:  end for
11:  for $t\in\text{shuffle}(\{T,\ldots,0\})$ do
12:   Compute critic loss:
13:    $\mathcal{L}_{\text{critic}}=\frac{1}{B}\sum_{i=1}^{B}\|V_{\phi}(x_{t}^{(i)},c^{(i)},t^{(i)})-r^{(i)}\|^{2}$
14:   Update critic: $\phi\leftarrow\phi-\alpha_{\text{critic}}\nabla_{\phi}\mathcal{L}_{\text{critic}}$
15:  end for
16:end forreturn $V_{\phi}$
B.2 Policy Training Algorithm
Algorithm 2 Policy Training with CGRU
1:Dataset of prompts $\mathcal{D}_{c}=\{c_{i}\}_{i=1}^{N}$, trained critic $V_{\phi}$, diffusion model $p_{\theta}$
2:Updated diffusion model $p_{\theta^{\prime}}$
3:Initialize policy parameters $\theta_{\text{old}}\leftarrow\theta$
4:Initialize empty trajectory buffer $\mathcal{T}=\emptyset$
5:for iteration =1=1 to II do
6:  for each prompt $c\sim\mathcal{D}_{c}$ do
7:   Generate trajectory: $x_{T}\sim\mathcal{N}(0,I)$, $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t},c)$ for $t=T,\ldots,1$
8:   Compute terminal reward: $r_{\text{final}}=r(x_{0},c)$
9:   Compute advantages: $A_{t}=r_{\text{final}}-V_{\phi}(x_{t},c,t)$ for all $t$
10:   Store trajectory $\tau=\{(x_{t},c,t,A_{t})\}_{t=1}^{T}$ in $\mathcal{T}$
11:  end for
12:  for $t\in\text{shuffle}(\{T,\ldots,0\})$ do
13:   Compute importance weights:
14:    $w^{(i)}=\frac{p_{\theta}(x_{t-1}^{(i)}|x_{t}^{(i)},c^{(i)})}{p_{\theta_{\text{old}}}(x_{t-1}^{(i)}|x_{t}^{(i)},c^{(i)})}$
15:   Compute policy update:
16:    $\nabla_{\theta}\mathcal{L}_{\text{policy}}=-\frac{1}{B}\sum_{i=1}^{B}w^{(i)}\log p_{\theta}(x_{t-1}^{(i)}|x_{t}^{(i)},c^{(i)})\cdot A_{t}^{(i)}$
17:   Update policy: $\theta\leftarrow\theta-\alpha_{\text{policy}}\nabla_{\theta}\mathcal{L}_{\text{policy}}$
18:  end for
19:end forreturn $p_{\theta}$
B.3 Complete CGRU Training Pipeline
Algorithm 3 Complete CGRU Training Pipeline
1:Concept to unlearn $\mathcal{C}$, base diffusion model $p_{\theta}^{\text{base}}$, prompt datasets $\mathcal{D}_{c}^{\text{forget}}$, $\mathcal{D}_{c}^{\text{retain}}$
2:Unlearned diffusion model $p_{\theta}^{\text{unlearned}}$
3:Step 1: Define Reward Function
4:Design $r(x_{0},c)$ to penalize concept $\mathcal{C}$ while preserving utility
5: (e.g., using CLIP-based similarity or classifier outputs)
6:Step 2: Train Critic
[…]
8:Step 3: Train Policy
[…]
Appendix C Feature-wise Linear Modulation (FiLM) Layers
C.1 FiLM Layer Formulation
C.2 Integration in Critic Architecture
C.3 Ablation Study: Impact of Timestep Awareness
To validate the effectiveness of our timestep-aware design, we conducted an ablation study comparing our architecture against a standard CLIP-based critic without timestep conditioning. We evaluated both models on their ability to predict the final reward from intermediate noisy latents.
Figure 4 presents the comparison results. The timestep-aware critic significantly outperforms the standard baseline, achieving an accuracy of 43.25% compared to 29.75%, and a macro precision of 58.52% compared to 34.47%. This substantial improvement confirms that conditioning the critic on the timestep via FiLM layers is essential for accurately estimating value functions from noisy states, particularly in the earlier stages of diffusion where the signal-to-noise ratio is low.
Figure 4: Ablation study comparing the performance of a standard critic vs. our timestep-aware critic on noisy latents. The timestep-aware model demonstrates significantly higher accuracy and macro precision, validating the importance of temporal conditioning for value estimation.
Appendix D Training Details
D.1 Hyperparameters
@@
Training epochs 100
@@
Training time $\sim$2 hours per run
D.2 Training Prompts
@@ D.4 Implementation Details
Appendix E Additional Experimental Results
E.1 Pareto Frontier Analysis
To better visualize the trade-off between Unlearning Accuracy (UA) and In-domain Retain Accuracy (IRA), we present a Pareto frontier plot in Figure 5. The plot compares CGRU against state-of-the-art baselines.
The green dashed line represents the baseline trade-off curve formed by methods such as SEOT, SPM, SHS, ESD, and UCE. Methods lying above this curve offer a superior balance between unlearning effectiveness and concept retention. As shown, CGRU (red star) is positioned well above this baseline frontier, indicating that it achieves significantly higher unlearning accuracy for its level of retention compared to the trend established by existing methods. This visualization confirms that our approach pushes the efficiency frontier of diffusion unlearning.
Figure 5: Trade-off between Unlearning Accuracy (UA) and In-domain Retain Accuracy (IRA). The green dashed line indicates the trend of existing baselines. CGRU (red star) lies above this trend, demonstrating a superior trade-off point favoring effective erasure while maintaining competitive retention.
E.2 Gradient Variance Analysis
E.3 Qualitative Progression Analysis
To provide deeper insight into the unlearning dynamics, we present qualitative examples showing the progression of concept erasure for three randomly selected classes. The prompts for these results were generated by ChatGPT using the following instruction: ”Generate 3 simple prompts for image generation for each of these 20 classes: {list of classes}”.
Figures 7, 8, and 9 illustrate the evolution of generated images throughout the training process. In each figure, the leftmost image corresponds to the untrained Stable Diffusion checkpoint (0 steps). Moving to the right, the images represent generations from checkpoints at increasing training steps. This progression highlights the gradual suppression of the target concept. We also observe that the convergence speed—the rate at which a concept is erased—varies across different classes, likely due to differences in concept frequency and semantic density within the pre-training data.
Figure 7: Qualitative progression for the ”Cats” class. The first image (left) is the untrained Stable Diffusion checkpoint, and images further to the right correspond to later training stages, showing the gradual erasure of the cat concept.
Figure 8: Qualitative progression for the ”Bears” class. The leftmost image represents the untrained baseline. As training progresses (moving right), the bear concept is effectively suppressed.
Figure 9: Qualitative progression for the ”Dogs” class. Progression from the untrained model (left) through the training process (right), demonstrating the removal of the dog concept.

### v3

@@ A.5 Advantage-weighted gradient form and Monte Carlo estimator
Since $B=0$, equation (10) reduces to the unbiased advantage-weighted gradient identity.
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{\tau\sim p_{\theta}(\cdot|c)}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)(r(x_{0},c)-b(x_{t},c,t))\right]$
@@
$\widehat{\nabla_{\theta}J}=\frac{1}{N}\sum_{n=1}^{N}\sum_{t=1}^{T}\left[\nabla_{\theta}\log p_{\theta}(x_{t-1}^{(n)}|x_{t}^{(n)},c^{(n)})(r(x_{0}^{(n)},c^{(n)})-b(x_{t}^{(n)},c^{(n)},t))\right]$
@@
Thus choosing $b(x_{t},c,t)=V_{\phi}(x_{t},c,t)$, precisely the value function in (8), and training $V_{\phi}$ to regress the observed terminal reward $r(x_{0},c)$ (Monte Carlo regression) approximates the optimal least-squares baseline $b^{\star}(x_{t})=\mathbb{E}[r(x_{0},c)|x_{t}]$ and thereby reduces variance of the estimator without introducing bias.
A.7 Off-policy extension via importance sampling
To enable multiple optimization steps on reused trajectories, we consider data collected under an earlier policy $p_{\theta_{\text{old}}}$. Correcting distribution shift with per-step importance weights yields the off-policy advantage-weighted estimator:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau\sim p_{\theta_{\text{old}}}}\left[\sum_{t=1}^{T}\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\theta_{\text{old}}}(x_{t-1}|x_{t},c)}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\,(r(x_{0},c)-b(x_{t},c,t))\right].$
This follows by standard importance sampling arguments Kakade and Langford (2002) applied to each term of the score-function expansion.
Appendix B Additional Experimental Results and Evaluation Protocol Details
B.1 UnlearnCanvas Evaluation Protocol Details
For each target concept, we evaluate the corresponding unlearned model on a fixed prompt grid combining content and style. Concretely, we use the template “A {class_name} image in {style} style”. For each of the 20 object classes and each of the 51 available styles (e.g., Watercolor, Cubism, Van_Gogh, Sketch, Neon_Lines), we generate 20 images, yielding $20\times 51\times 20$ samples per unlearned model. We run the UnlearnCanvas ViT classifier on each image and aggregate predicted class probabilities to compute per-class UA (for the target class) and IRA (mean over all non-target classes).
Table 4: Per-class metrics on UnlearnCanvas for CGRU. (higher is better).
Target class UA ($\uparrow$) IRA ($\uparrow$)
Architectures 94.50% 78.11%
Bears 97.94% 72.53%
Birds 94.29% 85.34%
Butterfly 96.46% 78.91%
Cats 95.33% 84.30%
Dogs 98.04% 75.64%
Fishes 97.25% 64.76%
Flame 98.43% 83.67%
Flowers 91.63% 89.78%
Frogs 96.08% 71.51%
Horses 98.71% 79.16%
Humans 95.28% 82.19%
Jellyfish 90.19% 81.19%
Rabbits 93.87% 96.42%
Sandwiches 92.58% 88.12%
Sea 95.98% 81.98%
Statues 96.22% 85.77%
Towers 94.27% 83.26%
Trees 93.51% 81.51%
Waterfalls 96.86% 88.75%
B.2 Misclassification matrices.
Figure 4: Per-class misclassification matrices (1/2). One matrix per target class.
Figure 5: Per-class misclassification matrices (2/2).
B.3 Additional Qualitative Results
We provide additional qualitative examples in the same format as Figure 3 in the main paper.
Figure 6: Qualitative progression examples (set 1).
Figure 7: Qualitative progression examples (set 2).
B.4 Gradient Variance Analysis
B.5 Importance Sampling Ratio Analysis
To assess the stability of trajectory reuse, we monitor the per-timestep importance ratio $\rho_{t}=p_{\theta}(x_{t-1}\mid x_{t},c)/p_{\theta_{\text{old}}}(x_{t-1}\mid x_{t},c)$ for each denoising step when updating the policy on trajectories collected by $\theta_{\text{old}}$. Figure 9 plots $\rho_{t}$ (aggregated over the batch and timesteps) across training before any clipping. We observe that the IS ratios remain tightly concentrated around $1$ with small variance and no heavy tails, indicating that policy updates stay in a local regime where $\theta$ does not drift far from $\theta_{\text{old}}$ between data collection and optimization. This empirically validates that multi-step reuse is effectively near on-policy in our setting and does not exhibit the instability typically associated with large IS weights; consequently, IS reweighting provides a safe correction for minor distribution mismatch while preserving stable training dynamics.
Figure 9: Importance-sampling (IS) ratio stability. IS ratios $\rho_{t}$ over training (before clipping), aggregated over timesteps and batch. Ratios remain tightly concentrated around 11, indicating stable near on-policy trajectory reuse.
Appendix C Training Details
C.1 Hyperparameters
@@
Training epochs 100–250 (class-dependent)
Gradient max-norm 1.0
Importance sampling ratio clipping $1-10^{-4},1+10^{-4}$]
@@
Training time $\sim$2–5 hours per run (class-dependent)
C.2 Training Prompts
@@ C.4 Implementation Details
Appendix D Training Algorithms
D.1 Critic Training Algorithm
Algorithm 1 Critic Training FOR CGRU
0: Dataset of prompts $\mathcal{D}_{c}=\{c_{i}\}_{i=1}^{N}$, diffusion model $p_{\theta}$, reward function $r(x_{0},c)$
0: Trained critic $V_{\phi}(x_{t},c,t)$
1: Initialize critic network $V_{\phi}$ with random parameters $\phi$
2: Initialize empty buffer $\mathcal{B}=\emptyset$
3:for epoch =1=1 to $E_{\text{critic}}$ do
4:  for each prompt $c\sim\mathcal{D}_{c}$ do
5:   Generate trajectory: $x_{T}\sim\mathcal{N}(0,I)$, $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t},c)$ FOR $t=T,\ldots,1$
6:   Compute terminal reward: $r_{\text{final}}=r(x_{0},c)$
7:   Store $(x_{t},c,t,r_{\text{final}})$ in $\mathcal{B}$ FOR all $t\in\{1,\ldots,T\}$
8:  end for
9:  for update =1=1 to $U_{\text{critic}}$ do
10:   Sample minibatch $\{(x_{t}^{(i)},c^{(i)},t^{(i)},r^{(i)})\}_{i=1}^{B}$ from $\mathcal{B}$
11:   Compute critic loss:
12:    $\mathcal{L}_{\text{critic}}=\frac{1}{B}\sum_{i=1}^{B}\|V_{\phi}(x_{t}^{(i)},c^{(i)},t^{(i)})-r^{(i)}\|^{2}$
13:   Update critic: $\phi\leftarrow\phi-\alpha_{\text{critic}}\nabla_{\phi}\mathcal{L}_{\text{critic}}$
14:  end for
15:end for
D.2 Policy Training Algorithm
Algorithm 2 Policy Training with CGRU
0: Dataset of prompts $\mathcal{D}_{c}=\{c_{i}\}_{i=1}^{N}$, critic $V_{\phi}$ (warm-started), diffusion model $p_{\theta}$
0: Updated diffusion model $p_{\theta^{\prime}}$
1: Initialize policy parameters $\theta_{\text{old}}\leftarrow\theta$
2: Initialize empty trajectory buffer $\mathcal{T}=\emptyset$
3: Initialize empty critic buffer $\mathcal{B}=\emptyset$
4:for iteration =1=1 to II do
5:  Set behavior policy parameters $\theta_{\text{old}}\leftarrow\theta$
6:  Clear trajectory buffer $\mathcal{T}\leftarrow\emptyset$
7:  Clear critic buffer $\mathcal{B}\leftarrow\emptyset$
8:  for each prompt $c\sim\mathcal{D}_{c}$ do
9:   Generate trajectory: $x_{T}\sim\mathcal{N}(0,I)$, $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t},c)$ FOR $t=T,\ldots,1$
10:   Compute terminal reward: $r_{\text{final}}=r(x_{0},c)$
11:   Compute advantages: $A_{t}=r_{\text{final}}-V_{\phi}(x_{t},c,t)$ FOR all $t$
12:   Store transitions $\tau=\{(x_{t},x_{t-1},c,t,A_{t})\}_{t=1}^{T}$ in $\mathcal{T}$
13:   Store $(x_{t},c,t,r_{\text{final}})$ in $\mathcal{B}$ FOR all $t\in\{1,\ldots,T\}$
14:  end for
15:  for update =1=1 to $U_{\text{critic}}^{\text{online}}$ do
16:   Sample minibatch $\{(x_{t}^{(i)},c^{(i)},t^{(i)},r^{(i)})\}_{i=1}^{B}$ from $\mathcal{B}$
17:   Compute critic loss:
18:    $\mathcal{L}_{\text{critic}}=\frac{1}{B}\sum_{i=1}^{B}\|V_{\phi}(x_{t}^{(i)},c^{(i)},t^{(i)})-r^{(i)}\|^{2}$
19:   Update critic: $\phi\leftarrow\phi-\alpha_{\text{critic}}\nabla_{\phi}\mathcal{L}_{\text{critic}}$
20:  end for
21:  for update =1=1 to $U_{\text{policy}}$ do
22:   Sample minibatch $\{(x_{t}^{(i)},x_{t-1}^{(i)},c^{(i)},t^{(i)},A_{t}^{(i)})\}_{i=1}^{B}$ from $\mathcal{T}$
23:   Compute importance weights:
24:    $w^{(i)}=\frac{p_{\theta}(x_{t-1}^{(i)}|x_{t}^{(i)},c^{(i)})}{p_{\theta_{\text{old}}}(x_{t-1}^{(i)}|x_{t}^{(i)},c^{(i)})}$
25:   Compute policy update:
26:    $\nabla_{\theta}\mathcal{L}_{\text{policy}}=-\frac{1}{B}\sum_{i=1}^{B}w^{(i)}\nabla_{\theta}\log p_{\theta}(x_{t-1}^{(i)}|x_{t}^{(i)},c^{(i)})\cdot A_{t}^{(i)}$
27:   Update policy: $\theta\leftarrow\theta-\alpha_{\text{policy}}\nabla_{\theta}\mathcal{L}_{\text{policy}}$
28:  end for
29:end for
D.3 Complete CGRU Training Pipeline
Algorithm 3 Complete CGRU Training Pipeline
0: Concept to unlearn $\mathcal{C}$, base diffusion model $p_{\theta}^{\text{base}}$, prompt datasets $\mathcal{D}_{c}$
0: Unlearned diffusion model $p_{\theta}^{\text{unlearned}}$
1:Step 1: Define Reward Function
2: Design $r(x_{0},c)$ to penalize concept $\mathcal{C}$ while preserving utility
3:  (e.g., using CLIP-based similarity or classifier outputs)
4:Step 2: Warm-start Critic
[…]
6:Step 3: Train Policy (with online critic updates)
[…]
Appendix E Feature-wise Linear Modulation (FiLM) Layers
E.1 FiLM Layer Formulation
E.2 Integration in Critic Architecture
E.3 Ablation Study
Figure 10 compares our timestep-aware critic against a standard CLIP-based critic without FiLM conditioning. Both models are evaluated on their ability to infer the concept signal used to define the terminal reward from intermediate noisy latents. Since the reward is computed from a 20-class concept classifier (via the target-class probability), we report classification metrics (accuracy and macro precision) on the induced discrete class predictions (argmax over the 20-way outputs). Test samples are drawn uniformly across timesteps to cover the full noise spectrum.
Figure 10: Ablation study comparing the performance of a standard critic vs. our timestep-aware critic on noisy latents. The timestep-aware critic achieves substantially higher accuracy (43.25% vs. 29.75%) and macro precision (58.52% vs. 34.47%).
@@ F.1 Fréchet Inception Distance (FID) Computation
F.2 CLIP Score Computation
For image–text alignment evaluation, we report the CLIP score as a cosine similarity between image and text embeddings, following standard practice. Specifically:
  • Model: We use the OpenAI CLIP ViT-L/14@336px encoder.
  • Computation: Given a generated image $I$ and its corresponding text prompt $T$, we obtain the normalized image embedding $f(I)$ and text embedding $g(T)$. The CLIP score is then
    $\text{CLIP}(I,T)=\frac{f(I)\cdot g(T)}{\|f(I)\|\,\|g(T)\|}\in[0,1].$
    Note: All CLIP scores reported in this paper are multiplied by 100 for readability.
  • Aggregation: For each model, we report the mean CLIP score over the entire evaluation set (30k prompts from MJHQ).
  • Variance reporting: We additionally compute the standard deviation across random seeds.
@@ Appendix G Large Language Model Usage
G.1 Safety-Aware Instruction for NSFW Prompt Generation
For the NSFW unlearning experiment (Section 5.4), we re-caption training images into text prompts using LLaVA-1.5-7B with the following safety-aware instruction:
CAPTION_USER_PROMPT = (
    "Describe this image in one concise sentence (15-25 words): "
    "main subject, pose/action, setting. "
    "Include what NSFW or suggestive elements are present in detail:
    nudity, revealing clothing, poses, etc."
    "Make it sound like a stable diffusion prompt."